OpenStudy (anonymous):

Express the function h(x) = (1/(x-2)) in the form f o g i.e. h(x) in form f(g(x)). If g(x) = x - 2, find the function f(x). Your answer is f(x) = ___________

OpenStudy (anonymous):

\[h(x) = 1\div(x-2) \] in the form fog i.e. h(x) in form f(g(x)). If g(x) = x-2, find the function f(x). Your answer is f(x) = _____

OpenStudy (anonymous):

here we have given that h(x)=1/(x-2) and g(x)=x-2 and h(x)=f(g(x)) so, we can rewrite it as...h(x)=f(g(x))=f((x-2)) as i plugged the value of g(x) given here.. here, we have given the function h(x) as 1/(x-2) and g(x) as x-2..that means our f(x) function must be 1/x...